choose the correct answerC-O in [Fe(CO)6]2+ have stretching bands higher in energy than those in free CO? Due to positively transition metal decreases
Answer #1

1. When a metal carbonyl has positive charge on it, the metal has deficiency of electrons due to which it shows the tendency to pull electrons. Also the tendency of metal to CO \pi * back bonding decreases. This shortens and strengthens the CO bond and hence the CO bond order and stretching frequency increases. Also the stretching bands have higher energy than in free CO.

So the correct option is D.

2. A. Total number of valence electrons = Mn (7) + \eta 5 C5H5 (5) + (CO)2 (2 x 2 = 4)

= 7 + 5 + 4 = 16

B. Total number of valence electrons = Cu (11) + \eta 5 C5H5 (5) + (CO)(2)

= 11 + 5 + 2 = 18

C. Total number of valence electrons = Cr (6) + \eta 5 C5H5 (5) + (CO)3 (2 x 3 = 6)

= 6 + 5 + 6 = 17

D. Total number of valence electrons = V (5) + \eta 5 C5H5 (5) + (CO)4 (2 x 4 = 8)

= 5 + 5 + 8 = 18

A and C do not follow 18 electron rule.